# On Edge Coloring of Multigraphs

### v4

@@ 5.2.1 A few basic properties
Let $\alpha$ and $\beta$ be colors in $\overline{\varphi}(T(y_{p-1}))$ with $v(\alpha)\prec_{\ell}v(\beta)$ and $\alpha\notin\varphi(T(v(\beta))-T_{n,q})$ ($T_{v(\beta)}-T_{n,q}=\emptyset$ if $v(\beta)\in T_{n,q}$). If $\alpha\in\overline{\varphi}(T_{n,q})-\{\gamma_{n}\}$, or $\alpha,\beta\notin D_{n,q}\cup\{\gamma_{n}\}$, or $\alpha,\beta$ are both closed for $T_{n,q}$, then $P_{v(\alpha)}(\alpha,\beta,\varphi)=P_{v(\beta)}(\alpha,\beta,\varphi)$. Additionally, if $\alpha\in\overline{\varphi}(T_{n,q})$ and $\alpha$ is closed for $T_{n,q}$, then $P_{v(\alpha)}(\alpha,\beta,\varphi)$ is the only $(\alpha,\beta)$-path intersecting $\partial(T_{n,q})$.
@@
Let $v(\alpha)=u$ and $v(\beta)=w$. We consider the following few cases.
Case I: $u,w\in T_{n,q}$.
If $T_{n,q}$ is closed for both $\alpha,\beta$, then no boundary edges of $T_{n,q}$ are colored by $\alpha$ or $\beta$. Thus we $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$, because $T_{n,q}$ is elementary by (5.0). Moreover, this is the only $(\alpha,\beta)$-path intersecting $T_{n,q}$, so Claim 5.6 holds in this case. Now we suppose $T_{n,q}$ is closed for $\alpha$ or $\beta$ but not both. We first consider the case $\alpha,\beta\neq\gamma_{n}$. By Claim 5.3 (could be applied on $\beta$ in place of $\alpha$ when $\beta$ is closed for $T_{n,q}$), there is at most one $(\alpha,\beta)$-path intersecting $T_{n,q}$. Thus $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$ is the unique $(\alpha,\beta)$-path intersecting $\partial(T_{n,q})$, as desired. We then assume $\beta=\gamma_{n}$. As $\gamma_{n}$ is closed for $T_{n,q}$ by Claim 5.3, $\alpha$ is not closed for $T_{n,q}$ by assumption. So we only need to show $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$. By Claim 5.1, $\alpha$ is closed for $T_{n,j}$, where $j$ is the smallest index with $\alpha\in\overline{\varphi}(T_{n,j})$. Thus both $\alpha$ and $\beta=\gamma_{n}$ are closed for $T_{n,j}$ by Claim 5.3, and therefore $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$ following (5.0). We now assume neither $\alpha$ nor $\beta$ is $T_{n,q}$-closed. As $\gamma_{n}$ is closed for $T_{n,q}$ by Claim 5.3, we have both $\alpha,\beta\neq\gamma_{n}$. In this case, we only need to show that $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$. By Claim 5.1, $\beta$ is closed for some $T_{n,j}$ with $\beta\in\overline{\varphi}(T_{n,j})$ and $0\leq j\leq q$. Because $u\prec w$, we have $u,w\in T_{n,j}$. Again by Claim 5.3, we can see that $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$ which is the unique $(\alpha,\beta)$-path intersecting $\partial(T_{n,j})$.
@@
In this case $\alpha\notin\varphi(T(w)-T_{n,q})$, $\beta\neq\gamma_{n}$, and $\beta$ is not closed for $T_{n,q}$. So we have $\alpha\neq\gamma_{n}$. We first assume that $\alpha$ is closed in $T_{n,q}$. By Claim 5.3, there is exactly one $(\alpha,\beta)$-path intersecting $T_{n,q}$, which is $P_{u}(\alpha,\beta,\varphi)$. Assume otherwise that $P_{u}(\alpha,\beta,\varphi)\neq P_{w}(\alpha,\beta,\varphi)$. Then $P_{w}(\alpha,\beta,\varphi)$ does not intersect $T_{n,q}$. Thus $\pi=\varphi/P_{w}(\alpha,\beta,\varphi)$ is $(T_{n,q},D_{n},\varphi)$-stable. Moreover, it is $T_{n}+f_{n}$ mod $\varphi$ by Claim 5.4, since $\alpha\neq\gamma_{n}$. Therefore, by Claim 5.2, $T_{n,q}$ is an ETT with a good hierarchy up to itself under $\pi$. Note that following TAA, the only case that $\beta\in\varphi(T(w)-T_{n,q})$ is $\beta=\delta_{n}$ and $q=0$. Thus because $\pi$ is $(T_{n,q},D_{n},\varphi)$-stable and $\alpha\notin\varphi(T(w)-T_{n,q})$, $\pi$ is also $(T(w),D_{n},\varphi)$-weakly stable. So $T(w)$ remains an ETT under $\pi$ by Lemma 3.1. Following Remark 3, $T(w)$ also has a good hierarchy under $\pi$. However, this a contradiction to (5.0) as $\alpha\in\overline{\pi}(w)\cap\overline{\pi}(T_{u})$.
Thus $\alpha$ is not closed in $T_{n,q}$. In this case we only need to prove $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$. Assume otherwise. Following Claim 5.1, we have $\alpha\notin\varphi(T_{n,q}-T_{n,r})$, where $r$ is the largest index with $r\leq q$ such that $\alpha$ is closed for $T_{n,r}$, and (1) $\alpha\in\cup_{\delta_{m}\in D_{n,k}}\Gamma^{k-1}_{m}$ for each $k$ with $r<k\leq q$. By the same argument as above for $T_{n,r}$ using Claim 5.3, we see that $P_{u}(\alpha,\beta,\varphi)$ is the only $(\alpha,\beta)$-path intersecting $T_{n,r}$, and $\pi=\varphi/P_{w}(\alpha,\beta,\varphi)$ is $(T_{n,r},D_{n},\varphi)$-stable. Similarly following Claim 5.4 and Claim 5.2, $T_{n,r}$ is an ETT with a good hierarchy up to itself under $\pi$. Since $\alpha\notin\varphi(T_{n,q}-T_{n,r})$ and $\beta\notin\overline{\varphi}(T(w)-w)$, we again see that $\pi$ is $(T(w),D_{n},\varphi)$-weakly stable. So the hierarchy of $T(w)$ satisfies Definition 4 (i), (ii) and (iii) under $\pi$ the same way as it is under $\varphi$. Following (1), we see that $T(w)$ also satisfies Definition 4 (iv), so $T(w)$ has a good hierarchy of $q$-levels. Again, this a contradiction to (5.0) as $\alpha\in\overline{\pi}(w)\cap\overline{\pi}(u)$.
@@
In this case, by (5.0), we have $\alpha\notin\varphi(T(w)-T_{n,q})$ and $\alpha,\beta\notin D_{n,q}\cup\overline{\varphi}(T_{n,q})$ So $\alpha,\beta\notin\varphi(T(w))$. Suppose on the contrary that $P_{u}(\alpha,\beta,\varphi)\neq P_{w}(\alpha,\beta,\varphi)$. Then it is easy to see that $\pi=\varphi/P_{w}(\alpha,\beta,\varphi)$ is $(T_{n,q},D_{n},\varphi)$-stable and $(T(w)-w,D_{n},\varphi)$-weakly stable. Moreover, $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi$ as under $\varphi$, where $h$ is the initial index of $\Theta_{n}$. So $\pi$ is $T_{n}+f_{n}$ mod $\varphi$. Thus following Lemma 3.1, Claim 5.2 and Remark 3, we see that $T(w)$ has a good hierarchy with $q$-levels, reaching a contradiction to (5.0) as $\alpha\in\overline{\pi}(w)\cap\overline{\pi}(T_{u})$. ∎
@@
    Let $P$ is an $(\alpha,\beta)$-path other than $P_{v(\alpha)}(\alpha,\beta,\varphi)$. If $\alpha\in\overline{\varphi}(T_{n,q})-\{\gamma_{n}\}$ and $\beta\neq\gamma_{n}$, or both $\alpha$ and $\beta$ are closed for $T_{n,q}$, then $\pi=\varphi/P$ is $T_{n}+f_{n}$ mod $\varphi$, $(T_{n,q},D_{n},\varphi)$-weakly stable, and $T_{n,q}$ has a good hierarchy of $(q-1)$-levels up to itself under $\pi$.
@@
    If $T_{n,q}\prec v(\alpha)\prec v(\beta)$, $\alpha\notin\varphi(T(v(\beta))-T_{n,q})$ and $\alpha,\beta\notin D_{n,q}$, then $\pi=\varphi/P$ is $(T_{n,q},D_{n},\varphi)$-stable and $T_{n}+f_{n}$ mod $\varphi$ for any $(\alpha,\beta)$-chain $P$. Moreover, $T$ is an ETT with a good hierarchy of $q$-levels under $\pi$.
@@
We first prove (i). If one of $\alpha$ and $\beta$ is closed in $T_{n,q}$, then $P$ does not intersect $T_{n,q}$ by Claim 5.6. So $\pi=\varphi/P$ is $(T_{n,q},D_{n},\varphi)$-stable. Moreover, it is $T_{n}$ mod $\varphi$ by Claim 5.4. So $T_{n,q}$ has a good hierarchy up to itself by Claim 5.2. We then assume neither $\alpha$ nor $\beta$ is closed in $T_{n,q}$. Because $\alpha\in\overline{\varphi}(T_{n,j})$, by Claim 5.1, there exist the largest index $r$ such that $\alpha$ or $\beta$ is closed in $T_{n,r}$. Thus by Claim 5.6 (could be applied on $\beta$ in place of $\alpha$ when $\beta$ is closed for $T_{n,r}$), $P$ does not intersect $T_{n,r}$, and $\pi=\varphi/P$ is $(T_{n,r},D_{n},\varphi)$-stable. Therefore, it is $T_{n}$ mod $\varphi$ by Claim 5.4, and $T_{n,r}$ has a good hierarchy up to itself by Claim 5.2. We claim that (1) $\alpha,\beta\notin\varphi(T_{n,q}-T_{n,r})$, unless $r=0$, $\beta=\delta_{n}$ with $\beta\notin\overline{\varphi}(T_{n,q})$, and $f_{n}$ is the only edge colored by $\beta$ in $E(T_{n,q}-T_{n,r})$ in this case. Indeed, if $\beta\in\overline{\varphi}(T_{n,q})$, then we have (1) immediately by Claim 5.1 by our choice of $r$. So $\beta\notin\overline{\varphi}(T_{n,q})$. So following Algorithm 3.1 and TAA, the only way that $\beta\in\varphi(T_{n,q}-T_{n,r})$ happens is $r=0$ and $\beta=\delta_{n}$ with $\beta\notin\overline{\varphi}(T_{n,q})$, and in this case $f_{n}$ is the only edge colored by $\beta$ in $E(T_{n,q}-T_{n,r})$. Note that we also have $\alpha\notin\varphi(T_{n,q}-T_{n,r})$ by Claim 5.1 with our choice of $r$. So (1) is established. Since $\pi$ is $(T_{n,r},D_{n},\varphi)$-stable, we see that $\pi$ is $(T_{n,q},D_{n},\varphi)$-weakly stable by (1). Because neither $\alpha$ nor $\beta$ is not closed for $T_{n,k}$ for each $r<k\leq q$ by our choice of $r$, it is then routine to check that $T_{n,q}$ indeed has a good hierarchy of $(q-1)$-levels up to itself under $\pi$. Now we prove (ii). Let $u=v(\alpha)$ and $w=v(\beta)$. So $\alpha\notin\varphi(T(w)-T_{n,q})$ and $\alpha,\beta\notin D_{n,q}\cup\overline{\varphi}(T_{n,q})$. Thus $\alpha,\beta\notin\varphi(T(w))$ and $\alpha,\beta\notin\Gamma^{q}$. By Claim 5.7, $P_{w}(\alpha,\beta,\varphi)=P_{u}(\alpha,\beta,\varphi)$. For any $(\alpha,\beta)$-chain $P$, it is easy to see that $\pi=\varphi/P$ is $(T_{n,q},D_{n},\varphi)$-stable and $(T(u)-u,D_{n},\varphi)$-weakly stable. Moreover, $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi$ as under $\varphi$, where $h$ is the initial index of $\Theta_{n}$. So $\pi$ is $T_{n}+f_{n}$ mod $\varphi$. Thus $T_{n,q}$ has a good hierarchy up to itself under $\pi$ by Claim 5.2. Now if $P\neq P_{u}(\alpha,\beta,\varphi)$, then $\pi$ is $(T,D_{n},\varphi)$-weakly stable, and therefore $T$ has a good hierarchy with $q$-levels following Remark 3. If $P=P_{u}(\alpha,\beta,\varphi)$, then since $\alpha,\beta\notin\varphi(T(w))$, $T$ can still be obtained from $T_{n,q}$ by TAA under $\pi$. Thus $T$ is an ETT under $\pi$. Because $\alpha,\beta\notin\Gamma^{q}$, even though colors of some edges in $T-T(w)$ may be switched between $\alpha$ and $\beta$, we still see that $T=T_{n,q+1}$ satisfies Definition 4 (i). So following Remark 3, $T$ remains an ETT with a good hierarchy of $q$-levels under $\pi$, as well. ∎

### v5

@@ 5.2.1 A few basic properties
Let $\alpha$ and $\beta$ be colors in $\overline{\varphi}(T(y_{p-1}))$ with $v(\alpha)\prec v(\beta)$ and $\alpha\notin\varphi(T(v(\beta))-T_{n,q})$ ($T_{v(\beta)}-T_{n,q}=\emptyset$ if $v(\beta)\in T_{n,q}$). If $\alpha\in\overline{\varphi}(T_{n,q})-\{\gamma_{n}\}$, or $\alpha,\beta\notin D_{n,q}\cup\{\gamma_{n}\}$, or $\alpha,\beta$ are both closed for $T_{n,q}$, then $P_{v(\alpha)}(\alpha,\beta,\varphi)=P_{v(\beta)}(\alpha,\beta,\varphi)$. Additionally, if $\alpha\in\overline{\varphi}(T_{n,q})$ and $\alpha$ is closed for $T_{n,q}$, then $P_{v(\alpha)}(\alpha,\beta,\varphi)$ is the only $(\alpha,\beta)$-path intersecting $\partial(T_{n,q})$.
@@
Let $v(\alpha)=u$ and $v(\beta)=w$. We consider the following few cases. Case I: $u,w\in T_{n,q}$.
If $T_{n,q}$ is closed for both $\alpha,\beta$, then no boundary edges of $T_{n,q}$ are colored by $\alpha$ or $\beta$. Thus we have $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$, because $T_{n,q}$ is elementary by (5.0). Moreover, this is the only $(\alpha,\beta)$-path intersecting $T_{n,q}$, so Claim 5.6 holds in this case.
Now we suppose $T_{n,q}$ is closed for $\alpha$ or $\beta$ but not both. We first consider the case $\alpha,\beta\neq\gamma_{n}$. By Claim 5.3 (could be applied on $\beta$ in place of $\alpha$ when $\beta$ is closed for $T_{n,q}$), there is at most one $(\alpha,\beta)$-path intersecting $T_{n,q}$. Thus $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$ is the unique $(\alpha,\beta)$-path intersecting $\partial(T_{n,q})$, as desired. We then assume $\beta=\gamma_{n}$. As $\gamma_{n}$ is closed for $T_{n,q}$ by Claim 5.3, $\alpha$ is not closed for $T_{n,q}$ by assumption. So we only need to show $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$. By Claim 5.1, $\alpha$ is closed for $T_{n,j}$, where $j$ is the smallest index with $\alpha\in\overline{\varphi}(T_{n,j})$. Thus both $\alpha$ and $\beta=\gamma_{n}$ are closed for $T_{n,j}$ by Claim 5.3, and therefore $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$ following (5.0).
We now assume neither $\alpha$ nor $\beta$ is $T_{n,q}$-closed. As $\gamma_{n}$ is closed for $T_{n,q}$ by Claim 5.3, we have both $\alpha,\beta\neq\gamma_{n}$. In this case, we only need to show that $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$. By Claim 5.1, $\beta$ is closed for some $T_{n,j}$ with $\beta\in\overline{\varphi}(T_{n,j})$ and $0\leq j\leq q$. Because $u\prec w$, we have $u,w\in T_{n,j}$. Again by Claim 5.3, we can see that $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$ which is the unique $(\alpha,\beta)$-path intersecting $\partial(T_{n,j})$.
@@
In this case $\alpha\notin\varphi(T(w)-T_{n,q})$, $\beta\neq\gamma_{n}$, and $\beta$ is not closed for $T_{n,q}$. So we have $\alpha\neq\gamma_{n}$. We first assume that $\alpha$ is closed in $T_{n,q}$. By Claim 5.3, there is exactly one $(\alpha,\beta)$-path intersecting $T_{n,q}$, which is $P_{u}(\alpha,\beta,\varphi)$. Assume otherwise that $P_{u}(\alpha,\beta,\varphi)\neq P_{w}(\alpha,\beta,\varphi)$. Then $P_{w}(\alpha,\beta,\varphi)$ does not intersect $T_{n,q}$. Thus $\pi=\varphi/P_{w}(\alpha,\beta,\varphi)$ is $(T_{n,q},D_{n},\varphi)$-stable. Moreover, it is $T_{n}+f_{n}$ mod $\varphi$ by Claim 5.4, since $\alpha\neq\gamma_{n}$. Therefore, by Claim 5.2, $T_{n,q}$ is an ETT with a good hierarchy up to itself under $\pi$. Note that following TAA, the only case that $\beta\in\varphi(T(w)-T_{n,q})$ is $\beta=\delta_{n}$ and $q=0$. Thus because $\pi$ is $(T_{n,q},D_{n},\varphi)$-stable and $\alpha\notin\varphi(T(w)-T_{n,q})$, $\pi$ is also $(T(w),D_{n},\varphi)$-weakly stable. So $T(w)$ remains an ETT under $\pi$ by Lemma 3.1. Following Remark 4, $T(w)$ also has a good hierarchy under $\pi$. However, this is a contradiction to (5.0) as $\alpha\in\overline{\pi}(w)\cap\overline{\pi}(T_{u})$.
Thus $\alpha$ is not closed in $T_{n,q}$. In this case we only need to prove $P_{u}(\alpha,\beta,\varphi)=P_{w}(\alpha,\beta,\varphi)$. Assume otherwise. Following Claim 5.1, we have $\alpha\notin\varphi(T_{n,q}-T_{n,r})$, where $r$ is the largest index with $r\leq q$ such that $\alpha$ is closed for $T_{n,r}$, and
(1) $\alpha\in\cup_{\delta_{m}\in D_{n,k}}\Gamma^{k-1}_{m}$ for each $k$ with $r<k\leq q$.
By the same argument as above for $T_{n,r}$ using Claim 5.3, we see that $P_{u}(\alpha,\beta,\varphi)$ is the only $(\alpha,\beta)$-path intersecting $T_{n,r}$, and $\pi=\varphi/P_{w}(\alpha,\beta,\varphi)$ is $(T_{n,r},D_{n},\varphi)$-stable. Similarly following Claim 5.4 and Claim 5.2, $T_{n,r}$ is an ETT with a good hierarchy up to itself under $\pi$. Since $\alpha\notin\varphi(T_{n,q}-T_{n,r})$ and $\beta\notin\overline{\varphi}(T(w)-w)$, we again see that $\pi$ is $(T(w),D_{n},\varphi)$-weakly stable. So the hierarchy of $T(w)$ satisfies Definition 4 (i), (ii) and (iii) under $\pi$ the same way as it is under $\varphi$. Following (1), we see that $T(w)$ also satisfies Definition 4 (iv), so $T(w)$ has a good hierarchy of $q$ levels. Again, this is a contradiction to (5.0) as $\alpha\in\overline{\pi}(w)\cap\overline{\pi}(u)$.
@@
In this case, by (5.0), we have $\alpha\notin\varphi(T(w)-T_{n,q})$ and $\alpha,\beta\notin D_{n,q}\cup\overline{\varphi}(T_{n,q})$ . So $\alpha,\beta\notin\varphi(T(w))$. Suppose on the contrary that $P_{u}(\alpha,\beta,\varphi)\neq P_{w}(\alpha,\beta,\varphi)$. Then it is easy to see that $\pi=\varphi/P_{w}(\alpha,\beta,\varphi)$ is $(T_{n,q},D_{n},\varphi)$-stable and $(T(w)-w,D_{n},\varphi)$-weakly stable. Moreover, $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi$ as under $\varphi$, where $h$ is the initial index of $\Theta_{n}$. So $\pi$ is $T_{n}+f_{n}$ mod $\varphi$. Thus following Lemma 3.1, Claim 5.2 and Remark 4, we see that $T(w)$ has a good hierarchy with $q$ levels, reaching a contradiction to (5.0) as $\alpha\in\overline{\pi}(w)\cap\overline{\pi}(T_{u})$. ∎
@@
    Let $P$ is an $(\alpha,\beta)$-path other than $P_{v(\alpha)}(\alpha,\beta,\varphi)$. If $\alpha\in\overline{\varphi}(T_{n,q})-\{\gamma_{n}\}$ and $\beta\neq\gamma_{n}$, or both $\alpha$ and $\beta$ are closed for $T_{n,q}$, then $\pi=\varphi/P$ is $T_{n}+f_{n}$ mod $\varphi$, $(T_{n,q},D_{n},\varphi)$-weakly stable, and $T_{n,q}$ has a good hierarchy of (q1)(q-1) levels up to itself under $\pi$.
@@
    If $T_{n,q}\prec v(\alpha)\prec v(\beta)$, $\alpha\notin\varphi(T(v(\beta))-T_{n,q})$ and $\alpha,\beta\notin D_{n,q}$, then $\pi=\varphi/P$ is $(T_{n,q},D_{n},\varphi)$-stable and $T_{n}+f_{n}$ mod $\varphi$ for any $(\alpha,\beta)$-chain $P$. Moreover, $T$ is an ETT with a good hierarchy of $q$ levels under $\pi$.
@@
We first prove (i). If one of $\alpha$ and $\beta$ is closed in $T_{n,q}$, then $P$ does not intersect $T_{n,q}$ by Claim 5.6. So $\pi=\varphi/P$ is $(T_{n,q},D_{n},\varphi)$-stable. Moreover, it is $T_{n}$ mod $\varphi$ by Claim 5.4. So $T_{n,q}$ has a good hierarchy up to itself by Claim 5.2.
We then assume neither $\alpha$ nor $\beta$ is closed in $T_{n,q}$. Because $\alpha\in\overline{\varphi}(T_{n,j})$, by Claim 5.1, there exist the largest index $r$ such that $\alpha$ or $\beta$ is closed in $T_{n,r}$. Thus by Claim 5.6 (could be applied on $\beta$ in place of $\alpha$ when $\beta$ is closed for $T_{n,r}$), $P$ does not intersect $T_{n,r}$, and $\pi=\varphi/P$ is $(T_{n,r},D_{n},\varphi)$-stable. Therefore, it is $T_{n}$ mod $\varphi$ by Claim 5.4, and $T_{n,r}$ has a good hierarchy up to itself by Claim 5.2. We claim that
(1) $\alpha,\beta\notin\varphi(T_{n,q}-T_{n,r})$, unless $r=0$, $\beta=\delta_{n}$ with $\beta\notin\overline{\varphi}(T_{n,q})$, and $f_{n}$ is the only edge colored by $\beta$ in $E(T_{n,q}-T_{n,r})$ in this case.
Indeed, if $\beta\in\overline{\varphi}(T_{n,q})$, then we have (1) immediately by Claim 5.1 by our choice of $r$. So $\beta\notin\overline{\varphi}(T_{n,q})$. So following Algorithm 3.1 and TAA, the only way that $\beta\in\varphi(T_{n,q}-T_{n,r})$ happens is $r=0$ and $\beta=\delta_{n}$ with $\beta\notin\overline{\varphi}(T_{n,q})$, and in this case $f_{n}$ is the only edge colored by $\beta$ in $E(T_{n,q}-T_{n,r})$. Note that we also have $\alpha\notin\varphi(T_{n,q}-T_{n,r})$ by Claim 5.1 with our choice of $r$. So (1) is established.
Since $\pi$ is $(T_{n,r},D_{n},\varphi)$-stable, we see that $\pi$ is $(T_{n,q},D_{n},\varphi)$-weakly stable by (1). Because neither $\alpha$ nor $\beta$ is closed for $T_{n,k}$ for each $r<k\leq q$ by our choice of $r$, it is then routine to check that $T_{n,q}$ indeed has a good hierarchy of (q1)(q-1) levels up to itself under $\pi$.
@@
Now we prove (ii). Let $u=v(\alpha)$ and $w=v(\beta)$. So $\alpha\notin\varphi(T(w)-T_{n,q})$ and $\alpha,\beta\notin D_{n,q}\cup\overline{\varphi}(T_{n,q})$. Thus $\alpha,\beta\notin\varphi(T(w))$ and $\alpha,\beta\notin\Gamma^{q}$. By Claim 5.7, $P_{w}(\alpha,\beta,\varphi)=P_{u}(\alpha,\beta,\varphi)$. For any $(\alpha,\beta)$-chain $P$, it is easy to see that $\pi=\varphi/P$ is $(T_{n,q},D_{n},\varphi)$-stable and $(T(u)-u,D_{n},\varphi)$-weakly stable. Moreover, $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\pi$ as under $\varphi$, where $h$ is the initial index of $\Theta_{n}$. So $\pi$ is $T_{n}+f_{n}$ mod $\varphi$. Thus $T_{n,q}$ has a good hierarchy up to itself under $\pi$ by Claim 5.2. Now if $P\neq P_{u}(\alpha,\beta,\varphi)$, then $\pi$ is $(T,D_{n},\varphi)$-weakly stable, and therefore $T$ has a good hierarchy with $q$ levels following Remark 4. If $P=P_{u}(\alpha,\beta,\varphi)$, then since $\alpha,\beta\notin\varphi(T(w))$, $T$ can still be obtained from $T_{n,q}$ by TAA under $\pi$. Thus $T$ is an ETT under $\pi$. Because $\alpha,\beta\notin\Gamma^{q}$, even though colors of some edges in TT(w)T-T(w) may be switched between $\alpha$ and $\beta$, we still see that $T=T_{n,q+1}$ satisfies Definition 4 (i). So following Remark 4, $T$ remains an ETT with a good hierarchy of $q$ levels under $\pi$, as well. ∎